# Which chromium-sulfur compounds exist as 2D material?

### v1

@@ 2.2 Absence of CrS2
Despite substantial efforts, the growth of CrS2 using our molecular beam epitaxy approach remained unachievable, both as mono- as well as multilayer. In all experiments, including those with variations in growth and annealing procedure, we always observed a plain hexagonal surface lattice with 0.341 nm as the lattice parameter. This rules out the formation of the predicted orthorhombic CrS2 phases 33, 52, the formation of the T’-phase 25, 31, and distorted T-phases 36, 31. H- and undistorted T-CrS2 display a hexagonal lattice, but with a significantly smaller lattice parameter of nm 31, 25, 34, 36.
@@
Growth variations involved lowering [see Figure S5(a)] and increasing the sulfur pressure during growth and annealing, as well as changing the growth temperature in the range from 160 K [see Figure S5(b)] to 600 K [see Figure S5(c)]. Particularly, we also changed from Gr/Ir(110) to Gr/Ir(111) as growth substrate [see Figure S6], without being able to detect an indication for CrS2 formation. Noteworthy, when grown on Gr/Ir(111) the apparent heights of the islands at V are 0.86 nm and 1.09 nm for Cr2S3-2D and CrS4-2D, respectively. This increase by about 0.07 nm compared to Gr/Ir(110) is primarily attributed to a larger van der Waals gap between CrxSy-2D and Gr. In fact, Gr is physisorbed on Ir(111) with weak chemical modulation 56, whereas it is chemisorbed on Ir(110) 54. This leads to stronger interactions with CrxSy-2D in the latter case and consequently to a reduced height above Gr.
@@
The MBE growth method used here has so far enabled us to reproduce whatever transition metal disulfide described in the literature that it was applied to 57, 58, 55. Therefore, the inability to grow CrS2-2D using the same method indicates that it may be worthwhile to reevaluate the evidence for the existence of CrS2.
@@
TEM images of CrS2 are provided in four publications 25, 26, 27, 28 and display a perfectly hexagonal lattice ruling out the observation of orthorhombic, distorted T- or T’-phases. Based on the specified net plane distances or scale bars provided, the lattice parameter is in the range of 0.33-0.36 nm, in descent agreement with the lattice parameter of 0.341 nm measured for Cr2S3-2D and CrS3-2D, but at variance with the calculated lattice parameters of nm 31, 25, 34, 36, 37 for the undistorted T- or H-phase. The measured CrS2 lattice parameters also agree with the reported lattice parameters of Cr2S3 thin films [see Figure 1(c)] ranging from 3.4 Å to 3.51 Å18, 22, 16, 17.
@@
The minimum heights of CrS2 flakes measured by atomic force microscopy are 2 nm 26, 1.5 nm 27, 0.75 nm 25, and 0.682 nm 28. Despite the three significant digits provided in the latter height, the noise level on each terrace is 0.2 nm. Given the experimental uncertainties, the measured heights are also compatible with Cr2S3-2D, or CrS3-2D and Cr2S3 thin films for the thicker layers.
@@
The characteristic Raman peaks of CrS2 scatter with a few wave numbers around 175 cm-1, 252 cm-1, 283 cm-1, and 359 cm-1 26, 27, 25, 28. It also has to be noted that Raman peaks are measured for thicker CrS2 films for signal reasons 27. Peak number and positions agree well with the ones reported for Cr2S3 thin films, specifically when considering their thickness-dependent variability 18, 19, 17.
@@
XPS characterization through the core level doublets S 2p3/2/S 2p1/2 and Cr 2p3/2/Cr 2p1/2 does neither provide a clear distinction between between CrS2 25, 26, 28and thin films of Cr2S3 18, 16, 20, 22, 17, 21. For instance, the data provided by Nair et al. 26 for CrS2 agree perfectly with the one provided by Cui et al. 21 for Cr2S3 thin films. Although XPS can be used for composition analysis, precise composition analysis is not feasible without proper calibration, taking depth-dependent attenuation of the signal into account and considering the depth distribution of the species in a layered material. It also has to be noted that the top three layers of S-Cr-S in thin Cr2S3 films and CrS2 are identical. Lastly, the prominent intensity peak at in X-ray diffraction of CrS2 25, 28, 27 is also present for thin Cr2S3 films 16, 20.
@@
Given the absence of CrS2 as a bulk material and the ambiguity surrounding the evidence for its existence as a 2D material, it may be valuable to consider how to obtain more conclusive evidence of CrS2.
2.3 Electronic structure from experiment
@@
Such a smeared-out behavior of the conduction band edge (CBE) has been observed in metallic MoS2 (see Figure S5 of the Supplementary Data of ref. 59). Detection of a sharp CBE for MoS2 in spectra was hampered by the fact that the conduction band minimum is located at the K-point of the first Brillouin zone (BZ). For such electrons with a large parallel momentum STS is relatively insensitive. As a result, precise energy assignment to the CBE is challenging. Here we assume a similar origin of the smeared-out CBE, which would imply a large parallel momentum of the electrons at the conduction band minimum. As an estimate, we suggest the CBE at V for both materials, yielding approximately band gaps of eV and eV for Cr2S3-2D and CrS4-2D, respectively. Furthermore, both spectra exhibit a pronounced dip at . Features at , including dips and peaks, are commonly observed in spectra of 2D materials 60 and are likely related to electronic correlations. However, the exact origin remains uncertain in this instance.
@@
Whether a 2D material with a full band gap is a semiconductor or a metal depends on the position of which may be affected by the substrate. In MoS2, a semiconductor-metal transition was driven by doping-induced work function lowering of the Gr substrate 59. Vacuum level alignment of Gr and MoS2 then implied a downshift of the MoS2 band structure until the CBE moved into the Fermi level and an interface dipole was formed by charge transfer from Gr to MoS2. We consider here for Cr2S3-2D and CrS4-2D a similar scenario. That is while the freestanding materials might be semiconducting, we speculate that a lower Gr work function implies, through vacuum level alignment for the CrxSy-2D compounds, could lead to a downward shift of their bands and, eventually, the formation of an interface dipole due to charge transfer into their conduction bands. Consequently, they become metallic. Such a scenario would imply that the work function of the CrxSy-2D compounds to be larger than the one of Gr. To substantiate this speculation, we estimated the work functions of the CrxSy-2D compounds using the method of field emission resonances 61, 62 and found them in fact to be larger by more than one electronvolt compared the one of Gr/Ir(110) (see Figure S7 of the SI).
@@
A notable feature in the spectrum of and CrS4-2D is a sharp rise peak at the VBE as highlighted by the inset in the lower left corner of Figure 6(b). Such a sharp feature might be related to a van Hove singularity of a surface state 63.
@@
Work function differences and the density of states also affect the measured apparent height in STM. The apparent heights of Cr2S3-2D and CrS4-2D were up to now only specified for V. This choice is meaningful because at positive bias the apparent heights are stable and vary only marginally with , as expected for a metal, for which the apparent height usually agrees rather well with the geometric height. Not unexpectedly, at negative bias in the range of the band gap, apparent heights up to 0.15 nm lower are measured (see Figure S8 in the SI).
@@
Finally, it is found that the electronic structure remains invariant with respect to the substrate, whether Gr/Ir(110) or Gr/Ir(111), as evidenced by comparison in Figure Figure S9. The invariance of the measured electronic structure is consistent with a weak interaction and the absence of significant hybridization with the two different substrates.
2.4 Density functional theory calculations
@@
Based on the experiments, a structure S-Cr-S-Cr-S consisting of five atomic planes is inferred for Cr2S3-2D, but the stacking sequence could not be established. In our DFT calculations, we tested the typical trigonal prismatic (1H) and octahedral (1T) coordination of Cr with respect to S, with stackings displaying aligned and shifted Cr positions between the two Cr planes (see Table S1 of the SI). To identify the most stable structure for Cr2S3-2D among the resulting eight possible crystal configurations, their total energies were calculated by DFT. The results, presented in Table S1 of the SI, give the lowest energy for the crystal structure with aligned 1T/1T Cr coordination, as illustrated in Figure 7(a) to (c). This crystal structure corresponds to the classical NiAs-structure and is the same as for bulk Cr2S3 and Cr5S6, when disregarding the presence of Cr vacancies in the bulk structures. The NiAs-structure of Cr2S3-2D is preferred by 0.17 eV per formula unit compared to a structure with the same octahedral coordination of the Cr atoms, but shifted Cr positions on the atomic planes, and by more than 0.6 eV compared to all other structures. The calculated lattice parameter of 0.344 nm agrees reasonably well with the experimental value of 0.341 nm. Furthermore, no superstructure is observed, consistent with the experimental results.
@@
Recalculating the band structure with does not change the overall picture, but the magnitude of the band gap is affected. It increases with increasing , reaches a maximum around 1.32 eV for eV and then drops to 0.84 eV for eV. Taking into account spin orbit coupling (SOC) decreases the bandgap size only slightly by 20 meV, for the complete list of bandgap sizes as a function of see Table S2 in the SI.
@@
A metallic state of Cr2S3-2D is obtained when the calculations are carried out considering the Gr substrate. Indeed, the CBE shifts below , indicating a semiconductor-metal transition of Cr2S3-2D (see projected density of states (PDOS) in Figure S10 in the SI).
@@
Next, we assess the magnetic properties of the Cr2S3-2D structure via DFT calculations. We find that the magnetic moment of the Cr+3 atoms (i.e., 3 ) is affected by SOC and reduced to 1.73 per Cr atom, indicating the presence of magnetism in the structure. To infer the magnetic ground state of Cr2S3-2D, intralayer Cr-Cr coupling was considered ferromagnetic (FM). This assumption is consistent with expectation based on the Goodenough-Kanamori-Anderson rule, which favors FM coupling when the bond angle Cr-S-Cr is close to 90° (in Cr2S3-2D it is 85.77°). For the interlayer coupling, we considered FM or A-type (or layerwise) antiferromagnetic (AFM) states, for a schematic representation, see Figure S11 in the SI. The energy difference between these two states is meV, that is the A-type AFM configuration is energetically preferred.
@@
The interlayer magnetic exchange splitting calculated by GGA is notably weak, emphasizing the importance of accurately accounting for electronic correlation effects when estimating magnetism. To address this, we analyzed the energetics of FM and AFM states by varying the parameters in GGA+U calculations. As the on-site Coulomb term increased, a transition from the AFM to the FM state was observed, as shown in Table S2 in the SI. This shift occurs because the super-exchange interlayer coupling is inversely related to the hopping parameter; thus, raising reduces the stability of AFM coupling relative to FM coupling.
@@
For CrS4-2D not only the stacking sequence, but also the position of the fractional Cr plane was uncertain from experimental data. Experimentally, LEED ruled out the fractional Cr plane to be the lowermost. Since for Cr2S3-2D octahedral (T) coordination of Cr atoms was preferred, we computed for CrS4-2D the four different structures consistent with these two constraints as outlined in Table S3 of the SI. We find the minimum energy configuration to be a NiAs-type structure with the fractional Cr plane in the center, as represented in Figures 8(a) to (c). This structure is preferred by 0.6 eV per formula unit. The DFT optimized configuration reflects the Cr5S6 motif of intermediate Cr layers filled only by [see Figures 1(c) and 1(e)]. The relaxed structure has a lattice constant of 0.347 nm, which agrees with an acceptable error to the experimental 0.341 nm. The thickness difference between the relaxed crystal structures of Cr2S3-2D and CrS4-2D is 0.26 nm, a value that closely aligns with the apparent height difference of 0.23 nm observed in STM.
@@
The ()-R30° superstructure visible in LEED and STM is according to our DFT calculations due to distortions caused by the missing Cr atoms in the middle Cr plane. The missing Cr atoms affect the position of S in the topmost plane. For three S atoms surrounding one Cr vacancy their displacements are highlighter by arrows in Figure 8(a). The DFT simulated constant-current STM topograph of CrS4-2D shown in the upper panel of Figure 8(d) closely resembles the experimental STM constant-current topograph shown in the lower panel.
@@
Band structure and PDOS of CrS4-2D are presented in Figure 8(e,f). The unfolded band structure has the conduction band minimum at the boundary of the original first BZ at the M-point, consistent with the smeared out CBE observed in the spectra in Figure 6(b) and as discussed above. The valence band maximum is located roughly in the middle between the M- and the -point. Thus, the electrons at the VBE have smaller parallel momentum, consistent with the sharp VBE seen in STS. The calculated band gap of CrS4-2D is with 0.33 eV clearly smaller than the one of Cr2S3-2D, consistent with experiment. However, a quantitative agreement of the band gap with the experiment is not expected. When is increased in the calculations, also the band gap increases and for a reasonable value of eV it amounts to 0.62 eV, close to the experimental result (see Table S4 of the SI).
@@
For CrS4-2D, the oxidation state of Cr atoms is also thanks to the identical stoichiometry of the 2D compound. The magnetic moment is retained and takes identical values to Cr2S3-2D, i.e., 3 without and 1.73 per Cr atom with SOC. To explore the magnetic ground state of CrS4-2D, we assumed similarly to Cr2S3-2D FM intralayer coupling. For the interlayer magnetic order three configurations were considered: one FM and two ferrimagnetic (FiM) ones. Two adjacent FM coupled Cr planes (one outer and the inner) couple AFM to the third (FiM-I), or the two outer Cr planes are both coupled AFM to the inner partial Cr plane (FiM-II), (see Figure S12). The relative energies between FM and FiM-I/FiM-II configurations are listed in Table S4. The results indicate that the ground state of CrS4-2D is FiM-I. CrS4-2D has an odd number of Cr planes and is a highly uncompensated antiferromagnet, as the magnetic moment of both outer Cr planes point in the same direction, while the fractional Cr plane contains magnetic moments pointing in the opposite direction. It results in ferrimagnet behavior displaying clearly spin-polarized bands, best visualized in Figure 8(e). Interestingly, CBE is dominated by spin-up states and VBE is composed of spin-down electrons. Like Cr2S3-2D, the interlayer magnetic coupling of CrS4-2D changes from AFM to FM as the is considered. The relative energies between FM and FiM-I/FiM-II configurations as a function of are given in Table S4 of the SI.
3 Conclusion

### v2

@@ II.2 Absence of CrS2
Despite substantial efforts, the growth of CrS2 using our molecular beam epitaxy approach remained unachievable, both as mono- as well as multilayer. In all experiments, including those with variations in growth and annealing procedure, we always observed a plain hexagonal surface lattice with 0.341 nm as the lattice parameter. This rules out the formation of the predicted orthorhombic CrS2 phases [33, 56], the formation of the T’-phase [25, 31], and distorted T-phases [36, 31]. H- and undistorted T-CrS2 display a hexagonal lattice, but with a significantly smaller lattice parameter of nm [31, 25, 34, 36].
@@
Growth variations involved lowering [see Figure S6(a), Supporting Information] and increasing the sulfur pressure during growth and annealing, as well as changing the growth temperature in the range from 160 K [see Figure S6(b), Supporting Information] to 600 K [see Figure S6(c), Supporting Information]. Particularly, we also changed from Gr/Ir(110) to Gr/Ir(111) as growth substrate [see Figure S7, Supporting Information], without being able to detect an indication for CrS2 formation. Noteworthy, when grown on Gr/Ir(111) the apparent heights of the islands at V are 0.86 nm and 1.09 nm for Cr2S3-2D and CrS4-2D, respectively. This increase by about 0.07 nm compared to Gr/Ir(110) is primarily attributed to a larger van der Waals gap between CrxSy-2D and Gr. In fact, Gr is physisorbed on Ir(111) with weak chemical modulation [61], whereas it is chemisorbed on Ir(110) [62]. This leads to stronger interactions with CrxSy-2D in the latter case and consequently to a reduced height above Gr.
The fact that the same phases of CrxSy-2D result on these two different substrates suggests that the outcome of growth does not depend on the substrate, be it metallic, insulating, or semiconducting, as long as it is sufficient inert to S and Cr and thus does not suppress the reaction of S and Cr with each other. Support for this view is from the fact single layer NbSe2 has the same structure and a nearly identical electronic structure, irrespective of whether grown on bi-layer Gr/SiC(0001), hexagonal boron nitride (h-BN) on Ir(111) or on bulk WSe2 [64]. Likewise, also MoS2 forms in the same way and under comparable conditions on Gr/Ir(111) and h-BN/Ir(111) [60].
@@
The MBE growth method used here has so far enabled us to reproduce whatever transition metal disulfide described in the literature that it was applied to [60, 65, 63]. Therefore, the inability to grow CrS2-2D using the same method indicates that it may be worthwhile to reevaluate the evidence for the existence of CrS2.
@@
TEM images of CrS2 are provided in four publications [25, 26, 27, 28] and display a perfectly hexagonal lattice ruling out the observation of orthorhombic, distorted T- or T’-phases. Based on the specified net plane distances or scale bars provided, the lattice parameter is in the range of 0.33-0.36 nm, in descent agreement with the lattice parameter of 0.341 nm measured for Cr2S3-2D and CrS3-2D, but at variance with the calculated lattice parameters of nm [31, 25, 34, 36, 37] for the undistorted T- or H-phase. The measured CrS2 lattice parameters also agree with the reported lattice parameters of Cr2S3 thin films [see Figure 1(c)] ranging from 3.4 Å to 3.51 Å[18, 22, 16, 17].
@@
The minimum heights of CrS2 flakes measured by atomic force microscopy are 2 nm [26], 1.5 nm [27], 0.75 nm [25], and 0.682 nm [28]. Despite the three significant digits provided in the latter height, the noise level on each terrace is 0.2 nm. Given the experimental uncertainties, the measured heights are also compatible with Cr2S3-2D, or CrS3-2D and Cr2S3 thin films for the thicker layers.
@@
The characteristic Raman peaks of CrS2 scatter with a few wave numbers around 175 cm-1, 252 cm-1, 283 cm-1, and 359 cm-1 [26, 27, 25, 28]. It also has to be noted that Raman peaks are measured for thicker CrS2 films for signal reasons [27]. Peak number and positions agree well with the ones reported for Cr2S3 thin films, specifically when considering their thickness-dependent variability [18, 19, 17].
@@
XPS characterization through the core level doublets S 2p3/2/S 2p1/2 and Cr 2p3/2/Cr 2p1/2 does neither provide a clear distinction between between CrS2 [25, 26, 28]and thin films of Cr2S3 [18, 16, 20, 22, 17, 21]. For instance, the data provided by Nair et al. [26] for CrS2 agree perfectly with the one provided by Cui et al. [21] for Cr2S3 thin films. Although XPS can be used for composition analysis, precise composition analysis is not feasible without proper calibration, taking depth-dependent attenuation of the signal into account and considering the depth distribution of the species in a layered material. It also has to be noted that the top three layers of S-Cr-S in thin Cr2S3 films and CrS2 are identical. Lastly, the prominent intensity peak at in X-ray diffraction of CrS2 [25, 28, 27] is also present for thin Cr2S3 films [16, 20].
@@
Given the absence of CrS2 as a bulk material and the ambiguity surrounding the experimental evidence for its existence as a 2D material, it may be valuable to consider how to obtain more conclusive experimental proof for it.
One might be tempted to use additional arguments for the existence of CrS2 as a 2D material from the analysis of the convex hull of the DFT-calculated formation energies in the Cr–S phase diagram. In fact the Computational 2D Materials Database C2DB [66, 67] finds the H-phase of single layer CrS2 on the convex hull, as do our own calculations discussed below (compare Figure 7). However, the relevance of this prediction is questioned by the fact that also the bulk compound of CrS2 is on the convex hull (see Materials Project database [68] and Figure 7), although its experimental synthesis has not been reported.
II.3 Electronic structure from experiment
@@
Such a smeared-out behavior of the conduction band edge (CBE) has been observed in metallic MoS2 (see Figure S5 of the Supplementary Data of ref. 69). Detection of a sharp CBE for MoS2 in spectra was hampered by the fact that the conduction band minimum is located at the K-point of the first Brillouin zone (BZ). For such electrons with a large parallel momentum STS is relatively insensitive. As a result, precise energy assignment to the CBE is challenging. Here we assume a similar origin of the smeared-out CBE, which would imply a large parallel momentum of the electrons at the conduction band minimum. As an estimate, we suggest the CBE at V for both materials, yielding approximately band gaps of eV and eV for Cr2S3-2D and CrS4-2D, respectively. Furthermore, both spectra exhibit a pronounced dip at . Features at , including dips and peaks, are commonly observed in spectra of 2D materials [70] and are likely related to electronic correlations. However, the exact origin remains uncertain in this instance.
@@
Whether a 2D material with a full band gap is a semiconductor or a metal depends on the position of which may be affected by the substrate. In MoS2, a semiconductor-metal transition was driven by doping-induced work function lowering of the Gr substrate [69]. Vacuum level alignment of Gr and MoS2 then implied a downshift of the MoS2 band structure until the CBE moved into the Fermi level and an interface dipole was formed by charge transfer from Gr to MoS2. We consider here for Cr2S3-2D and CrS4-2D a similar scenario. That is while the freestanding materials might be semiconducting, we speculate that a lower Gr work function implies, through vacuum level alignment for the CrxSy-2D compounds, could lead to a downward shift of their bands and, eventually, the formation of an interface dipole due to charge transfer into their conduction bands. Consequently, they become metallic. Such a scenario would imply that the work function of the CrxSy-2D compounds to be larger than the one of Gr. To substantiate this speculation, we estimated the work functions of the CrxSy-2D compounds using the method of field emission resonances [71, 72] and found them in fact to be larger by more than one electronvolt compared the one of Gr/Ir(110) (see Figure S8, Supporting Information).
@@
A notable feature in the spectrum of and CrS4-2D is a sharp rise peak at the VBE as highlighted by the inset in the lower left corner of Figure 6(b). Such a sharp feature might be related to a van Hove singularity of a surface state [73].
@@
Work function differences and the density of states also affect the measured apparent height in STM. The apparent heights of Cr2S3-2D and CrS4-2D were up to now only specified for V. This choice is meaningful because at positive bias the apparent heights are stable and vary only marginally with , as expected for a metal, for which the apparent height usually agrees rather well with the geometric height. Not unexpectedly, at negative bias in the range of the band gap, apparent heights up to 0.15 nm lower are measured (see Figure S9, Supporting Information).
@@
Finally, it is found that the electronic structure remains invariant with respect to the substrate, whether Gr/Ir(110) or Gr/Ir(111), as evidenced by comparison in Figure Figure S10 (Supporting Information). The invariance of the measured electronic structure is consistent with a weak interaction and the absence of significant hybridization with the two different substrates.
II.4 Density functional theory calculations
@@
Based on the experiments, a structure S-Cr-S-Cr-S consisting of five atomic planes is inferred for Cr2S3-2D, but the stacking sequence could not be established. In our DFT calculations, we tested the typical trigonal prismatic (1H) and octahedral (1T) coordination of Cr with respect to S, with stackings displaying aligned and shifted Cr positions between the two Cr planes (see Table S1, Supporting Information). To identify the most stable structure for Cr2S3-2D among the resulting eight possible crystal configurations, their total energies were calculated by DFT. The results, presented in Table S1 (Supporting Information), give the lowest energy for the crystal structure with aligned 1T/1T Cr coordination, as illustrated in Figure 8(a) to (c). This crystal structure corresponds to the classical NiAs-structure and is the same as for bulk Cr2S3 and Cr5S6, when disregarding the presence of Cr vacancies in the bulk structures. The NiAs-structure of Cr2S3-2D is preferred by 0.17 eV per formula unit compared to a structure with the same octahedral coordination of the Cr atoms, but shifted Cr positions on the atomic planes, and by more than 0.6 eV compared to all other structures. The calculated lattice parameter of 0.344 nm agrees reasonably well with the experimental value of 0.341 nm. Furthermore, no superstructure is observed, consistent with the experimental results.
In the convex hull plot of Figure 7, the formation energy of the Cr2S3-2D is seen to be just disfavored by about 20 meV/atom to the bulk phase with the same stoichiometry (but different structure). The DFT-calculated van der Waals binding energy of Cr2S3-2D to Gr of 36 meV/atom (17.5 meV/Å2) is larger than the lack in formation energy and thus favors the formation of the 2D phase over the bulk phase with the same composition. However, we note that these energy differences are not significantly outside the error margin of DFT calculations. Additionally, the thermal stability of the new 2D-phases was also confirmed through molecular dynamics simulations conducted at 700 K for 10 ps (see video files as Supporting Information). The absence of imaginary frequencies in the DFT calculated phonon dispersion, shown in Figure 2 of ref. 40, is also in line with the experimentally observed stability of Cr2S3-2D.
A noteworthy feature of the DFT-optimized Cr2S3-2D structure is the asymmetry of the Cr atomic planes with respect to the S ones. As well visible in the side view ball models of the final DFT geometry in Figures 8(b) and (c), the Cr planes are relaxed outwards by about nm away from the center position between the S planes. Consequently the bond length of the surface S atoms of 0.233 nm is considerably shorter than the bond length of the center S atoms of length 0.253 nm. The bond shortening nm of the surface S atoms is interpreted as a bond strengthening in consequence of their lower coordination. In fact, an energy gain of 0.26 eV per formula unit is associated with the relaxation. The energy gain due to bond strengthening of the surface S atoms may explain in part, why the formation energies of the 2D compounds are only higher by about 20 meV/atom as compared to a bulk phase with the same composition.
@@
Recalculating the band structure with does not change the overall picture, but the magnitude of the band gap is affected. It increases with increasing , reaches a maximum around 1.32 eV for eV and then drops to 0.84 eV for eV. Taking into account spin orbit coupling (SOC) decreases the bandgap size only slightly by 20 meV, for the complete list of bandgap sizes as a function of see Table S2 (Supporting Information).
Given the semiconducting nature of free-standing Cr2S3-2D, one can estimate the associated exciton binding energy. Using Green’s function–Bethe–Salpeter equation, we calculated this value to be of the order of 0.38 eV, see Figure S11 (Supporting Information). The exciton binding energy for Cr2S3-2D lies between the reported values for CrSBr bulk (0.25 eV) [74] and monolayer (0.5 eV) [75].
@@
A metallic state of Cr2S3-2D is obtained when the calculations are carried out considering the Gr substrate. Indeed, the CBE shifts below , indicating a semiconductor-metal transition of Cr2S3-2D (see projected density of states (PDOS) in Figure S12, Supporting Information).
@@
Next, we assess the magnetic properties of the Cr2S3-2D structure via DFT calculations. We find that the magnetic moment of the Cr+3 atoms (i.e., 3 ) is affected by SOC and reduced to 1.73 per Cr atom, indicating the presence of magnetism in the structure. To infer the magnetic ground state of Cr2S3-2D, intralayer Cr-Cr coupling was considered ferromagnetic (FM). This assumption is consistent with expectation based on the Goodenough-Kanamori-Anderson rule, which favors FM coupling when the bond angle Cr-S-Cr is close to 90° (in Cr2S3-2D it is 85.77°). For the interlayer coupling, we considered FM or A-type (or layerwise) antiferromagnetic (AFM) states, for a schematic representation, see Figure S13 (Supporting Information). The energy difference between these two states is meV, that is the A-type AFM configuration is energetically preferred.
@@
The interlayer magnetic exchange splitting calculated by GGA is notably weak, emphasizing the importance of accurately accounting for electronic correlation effects when estimating magnetism. To address this, we analyzed the energetics of FM and AFM states by varying the parameters in GGA+U calculations. As the on-site Coulomb term increased, a transition from the AFM to the FM state was observed, as shown in Table S2 (Supporting Information). This shift occurs because the super-exchange interlayer coupling is inversely related to the hopping parameter; thus, raising reduces the stability of AFM coupling relative to FM coupling.
@@
For CrS4-2D not only the stacking sequence, but also the position of the fractional Cr plane was uncertain from experimental data. Experimentally, LEED ruled out the fractional Cr plane to be the lowermost. Since for Cr2S3-2D octahedral (T) coordination of Cr atoms was preferred, we computed for CrS4-2D the four different structures consistent with these two constraints as outlined in Table S3 (Supporting Information). We find the minimum energy configuration to be a NiAs-type structure with the fractional Cr plane in the center, as represented in Figures 9(a) to (c). This structure is preferred by 0.6 eV per formula unit. The DFT optimized configuration reflects the Cr5S6 motif of intermediate Cr layers filled only by [see Figures 1(c) and 1(e)]. The relaxed structure has a lattice constant of 0.347 nm, which agrees with an acceptable error to the experimental 0.341 nm. The thickness difference between the relaxed crystal structures of Cr2S3-2D and CrS4-2D is 0.26 nm, a value that aligns closely with the apparent height difference of 0.23 nm observed in STM.
Similar to Cr2S3-2D, for CrS4-2D the gain in van der Waals binding of 33 meV/atom (19.8 meV/Å2) outweighs the lack in binding energy of about 21 meV/atom compared to a bulk crystal with the same composition. As for Cr2S3-2D, the thermal stability of CrS4-2D was confirmed through molecular dynamics simulations (see video file as Supporting Information). Lastly, also for CrS4-2D clear surface relaxation effects are visible. Again the outer Cr planes are relaxed outwards towards the surface by nm, making the S surface bonds shorter by nm.
@@
The ()-R30° superstructure visible in LEED and STM is according to our DFT calculations due to distortions caused by the missing Cr atoms in the middle Cr plane. The missing Cr atoms affect the position of S in the topmost plane. For three S atoms surrounding one Cr vacancy, their displacements are highlighted by arrows in Figure 9(a). The DFT simulated constant-current STM topograph of CrS4-2D shown in the upper panel of Figure 9(d) closely resembles the experimental STM constant-current topograph shown in the lower panel.
@@
Band structure and PDOS of CrS4-2D are presented in Figure 9(e,f). The unfolded band structure has the conduction band minimum at the boundary of the original first BZ at the M-point, consistent with the smeared out CBE observed in the spectra in Figure 6(b) and as discussed above. The valence band maximum is located roughly in the middle between the M- and the -point. Thus, the electrons at the VBE have smaller parallel momentum, consistent with the sharp VBE seen in STS. The calculated band gap of CrS4-2D is with 0.33 eV clearly smaller than the one of Cr2S3-2D, consistent with experiment. However, a quantitative agreement of the band gap with the experiment is not expected. When is increased in the calculations, also the band gap increases and for a reasonable value of eV it amounts to 0.62 eV, close to the experimental result (see Table S4, Supporting Information).
@@
For CrS4-2D, the oxidation state of Cr atoms is also thanks to the identical stoichiometry of the 2D compound. The magnetic moment is retained and takes identical values to Cr2S3-2D, i.e., 3 without and 1.73 per Cr atom with SOC. To explore the magnetic ground state of CrS4-2D, we assumed similarly to Cr2S3-2D FM intralayer coupling. For the interlayer magnetic order three configurations were considered: one FM and two ferrimagnetic (FiM) ones. Two adjacent FM coupled Cr planes (one outer and the inner) couple AFM to the third (FiM-I), or the two outer Cr planes are both coupled AFM to the inner partial Cr plane (FiM-II), (see Figure S14, Supporting Information). The relative energies between FM and FiM-I/FiM-II configurations are listed in Table S4 (Supporting Information). The results indicate that the ground state of CrS4-2D is FiM-I. CrS4-2D has an odd number of Cr planes and is a highly uncompensated antiferromagnet, as the magnetic moment of both outer Cr planes point in the same direction, while the fractional Cr plane contains magnetic moments pointing in the opposite direction. It results in ferrimagnet behavior displaying clearly spin-polarized bands, best visualized in Figure 9(e). Interestingly, CBE is dominated by spin-up states and VBE is composed of spin-down electrons. Like Cr2S3-2D, the interlayer magnetic coupling of CrS4-2D changes from AFM to FM as the is considered. The relative energies between FM and FiM-I/FiM-II configurations as a function of are given in Table S4 (Supporting Information).
III Conclusion